i am asking for help for a worded quadratic equation.the lenght of a rectangle is 5cm more than its width and the areais 24cm square find the length and the width of triangle.i really need help with this equation i am wondering if i could get the working out thank u
Answer by scott8148(6628)

let x=width, so x+5=length ... x(x+5)=24

x^2+5x-24=0 ... factoring gives (x+8)(x-3)=0